# See or Say Graphs: Agent-Driven Scalable Graph Understanding with Vision-Language Models

### v1

@@ Abstract
See or Say Graphs: Agent-Driven Scalable Graph Understanding
with Vision-Language Models
@@
**footnotetext: These authors contributed equally to this work.footnotetext: Corresponding author.
@@ 1 Introduction
Recently, vision-language models (VLMs) have shown potential for general-purpose understanding of graph data Tang et al. (2024a); Kong et al. (2025b); Ding et al. (2025b), offering new paradigms for solving real-world problems naturally represented as graphs. A central research focus of this emerging field is to enhance the foundational capability of VLMs to understand graph structures Ren et al. (2024), particularly for large-scale graphs, without relying on external tools such as code execution or software systems Ding et al. (2025a); Feng et al. (2025). To this end, existing studies have explored two primary graph input modalities for VLMs: textual graph descriptions [^1] Many early studies on text-modality-based graph understanding were developed on LLMs. As VLMs extend LLMs with visual capabilities, this paper presents a unified discussion from the perspective of graph understanding with VLMs., which convert structural information into natural language for indirect memory and reasoning over node properties and relations Jiang et al. (2023); Jin et al. (2024); Zhang et al. (2024a); and visual graph representations, which render nodes and edges as structured images to support direct structural perception Li et al. (2024c); Zhu et al. (2025). More recent efforts further combine these modalities to improve structural understanding and complex reasoning on graphs Wei et al. (2024).
@@
Figure 1: Comparison of textual and visual modalities across different types of graph understanding tasks
@@
However, despite these advances, current VLMs still lack a unified framework that systematically addresses the scalability challenges and the difficulties in coordinating textual and visual modalities inherent in graph understanding. Specifically:
@@
Scalability remains a fundamental limitation for both textual and visual graph inputs in VLMs. Given their limited input token capacities, textual graph descriptions are constrained by sequence length, making it difficult to encode the full structure of large graphs Chen et al. (2024a); Yuan et al. (2025a). Visual graph inputs, on the other hand, are restricted by image resolution, resulting in the loss of structural details as graph size increases Zhu et al. (2025).
@@
In addition, VLMs face the lack of an effective mechanism to coordinate different input modalities for graph understanding, which prevents them from fully leveraging the distinct strengths and complementary advantages of textual and visual modalities. As illustrated in Figure 1, under the token length limits of VLMs, the text modality is more effective for simple property tasks such as counting nodes, measuring degrees, and assessing connectivity. In these cases, the information can be directly located and reasoned over with minimal effort. By contrast, the visual modality often performs worse because the limited resolution of global visual graphs blurs structural details, causing VLMs to misidentify relevant elements and leading to performance degradation.
@@
For tasks that require local and structurally complex reasoning, such as shortest-path computation or cycle detection, high-resolution local visual subgraphs offer a clear advantage for VLMs. They provide a “what-you-see-is-what-you-get” view of the topology and naturally support stepwise reasoning. In contrast, textual descriptions are often lengthy and lack structural clarity, leaving VLMs with only indirect language cues and thus little structural grounding for effective complex graph reasoning.
Therefore, a natural conclusion is that realizing the full potential of VLMs in graph understanding requires addressing these two challenges.
@@
This entails, on the one hand, preserving as much task-relevant graph information as possible within limited input tokens, and on the other hand, assigning tasks according to modality strengths: letting the text modality efficiently address simple property tasks, while leveraging the visual modality to directly support local and structurally complex tasks.
@@
To this end, we propose GraphVista, a unified framework for graph understanding with VLMs. For the scalability challenges, GraphVista is inspired by the retrieval-augmented generation (RAG) paradigm Lewis et al. (2020). It organizes graph information hierarchically according to structural importance and stores multi-granularity textual descriptions in a lightweight GraphRAG base. For a given graph understanding task, GraphVista retrieves only the relevant information before input, thereby compressing redundant context and preserving key reasoning elements. When necessary, it can also generate high-resolution visual subgraphs from the retrieved local information to support complex reasoning tasks.
@@
To enable effective coordination across modalities, GraphVista introduces a planning agent that parses graph understanding tasks and routes them to the appropriate modality. The textual modality branch, enhanced with the RAG techniques, efficiently handles simple property tasks by retrieving and reasoning over concise textual descriptions. In contrast, the visual modality branch employs the Visual Graph Thoughts Agent, which constructs a multimodal chain of thought called Visual Graph Thoughts over high-resolution visual subgraphs, enabling VLMs to perform stepwise visual reasoning grounded in explicit structural evidence for structurally complex tasks.
The main contributions of this paper are listed as follows:
@@
    We present the first systematic analysis of the scalability challenges and coordination gaps that hinder VLMs in graph understanding, providing both conceptual insights and empirical evidence.
@@
    We propose GraphVista, the first unified framework for graph understanding with VLMs, designed around two objectives: improving scalability and enabling coordination across modalities. To this end, GraphVista integrates a set of innovative techniques, such as a hierarchical GraphRAG base, a planning agent for task routing, and a Multimodal Graph Thoughts agent for complex reasoning.
@@
    We introduce Grena, a scalable benchmark that supports step-level training of Visual Graph Thoughts and evaluates graph understanding from simple property to complex local tasks.
@@
    Extensive experiments across diverse graph understanding tasks demonstrate the effectiveness of GraphVista, highlighting its ability to scale to large graphs and to exploit the complementary strengths of textual and visual modalities.
@@ 2 Related Work
2.1 Text Modality-based Graph Understanding
Most existing studies convert graphs into textual descriptions to enable VLM-based graph understanding, primarily focusing on benchmark construction and capability analysis. Recent benchmarks evaluate VLMs on basic graph element comprehension Wang et al. (2023); Guo et al. (2023); Wu et al. (2024) and graph-theoretic reasoning tasks Chen et al. (2024a); Li et al. (2024b); Tang et al. (2024b); Luo et al. (2024); Yuan et al. (2025a), indicating that the research in this direction remains in an early exploratory phase. Methodologically, Ge et al. (2025); Wang et al. (2025) studied the influence of graph description order, while Cao et al. (2025) identified the “Lost in the Middle” issue in graph sequence. Building upon these findings, Guan et al. (2025) analyzed VLM attention patterns over graph data, and Peng et al. (2025); Zhang et al. (2025) explored task transferability to broader reasoning domains. However, these strategies still face scalability bottlenecks and struggle with complex reasoning tasks.
@@ 2.1 Text Modality-based Graph Understanding
2.2 Visual Modality-based Graph Understanding
Visual modality-based graph understanding is still in its infancy. VisionGraph Li et al. (2024c) and VGCure Zhu et al. (2025) benchmark VLMs on graph perception and reasoning, revealing difficulties in capturing structural information. Wei et al. (2024) further introduces a hybrid multimodal representation directly compatible with VLMs. However, these studies remain empirical and fail to address the scalability, structural understanding, and reasoning limitations inherent to vision-only representations.
@@ 3.1 Overview
In this section, we introduce GraphVista, an agent-driven framework designed to address the challenges of scalability and modality coordination in graph understanding, as shown in Figure 2.
GraphVista operates under the coordination of the planning agent, which orchestrates the entire reasoning workflow. Initially, the planning agent constructs a lightweight Hierarchical GraphRAG Base $\mathcal{K}$ by partitioning the graph based on topological centrality. Upon receiving a question, it parses the task and routes it to the appropriate reasoning modality. For Simple Property Tasks, the text modality-based branch is activated, employing GraphRAG-based retrieval to extract relevant context from $\mathcal{K}$. For Complex Local Tasks that require fine-grained structural reasoning, the task is delegated to the visual modality-based branch. The Visual Graph Thoughts Agent extracts visual subgraphs from $\mathcal{K}$ and performs iterative multimodal reasoning grounded in visual evidence.
The rest of this section details our methodology. We begin by introducing the construction of $\mathcal{K}$ in §3.2. Next, we describe the planning agent-based task routing mechanism in §3.3. Finally, we elaborate on the visual modality-based and text modality-based graph understanding modules in §3.4 and §3.5, respectively.
@@ 3.2 Hierarchical GraphRAG Base
A primary challenge for LLMs and VLMs in understanding graph structures is their limited input capacity. This constraint makes it infeasible to encode an entire graph, often causing a loss of critical structural information. To address this limitation, we draw inspiration from RAG and propose building $\mathcal{K}$. Unlike conventional RAG pipelines Guo et al. (2025); Jin et al. (2025) that rely on external knowledge sources, our approach constructs a compact multi-level structural representation directly from the graph and stores it in a lightweight GraphRAG Base, enabling scalable reasoning without external dependencies [^2] For small graphs with $|V|\leq 15$, we skip storage and feed the entire graph directly into the VLMs..
We partition the graph $G=(V,E)$ with $|V|$ nodes into three tiers based on the topological centrality of its nodes. Each tier uses a different storage granularity, significantly reducing storage needs while preserving the structural details essential for downstream tasks.
@@
This tier consists of nodes with the highest structural influence, functioning as hubs that govern information flow and shape global graph properties. We employ PageRank, a widely adopted centrality measure that reflects both connectivity and reachability, to identify the top $K_{1}\%$ of nodes as the core nodes. For each core node $v$, its complete 2-hop neighborhood $\mathcal{N}_{2}(v)$ is preserved to capture fine-grained local structural information.
@@
This tier contains nodes located along intercommunity paths, which are essential for preserving global connectivity. These nodes are selected using betweenness centrality, which quantifies how frequently a node lies on shortest paths between other nodes. The next $K_{2}\%$ of nodes are retained, and for each backbone node $v$, only its 1-hop neighborhood $\mathcal{N}_{1}(v)$ is stored to balance representational fidelity with storage efficiency.
@@ 3.3 Planning Agent-based Task Routing
This module employs a planning agent to allocate each task to the optimal reasoning modality, either textual or visual. The routing procedure comprises two stages. In the Task Parsing stage, the agent analyzes the natural language question to extract the task intent and key entities. In the Task Categorization stage, the parsed task is classified and dispatched to the corresponding reasoning module.
Task Parsing
Given a natural language question $Q$ over $G$, the planning agent first performs semantic parsing to identify the task type $T$ and extract the set of key entities $\mathcal{E}$:
@@
, where $f_{\text{parse}}$ denotes a semantic parsing function implemented by the planning agent. The agent applies few-shot prompting to infer task semantics, producing a task type $T$ (e.g., neighbor identification) and a set of key entity pairs $\mathcal{E}={(v_{i},v_{j})\mid v_{i},v_{j}\in V}$ referenced in the question. The task type $T$ determines the routing strategy, while the entity set $\mathcal{E}$ provides essential context for downstream modules. To ensure uniform processing, queries involving a single entity (e.g., “What is the degree of node $v_{i}$?”) are reformulated as $(v_{i},\emptyset)$ by introducing a null entity placeholder.
Task Categorization and Execution Planning
Following parsing, the planning agent routes the task to the appropriate module based on $T$ and formulates a high-level execution plan [^4] The planning agent first matches the question against predefined task templates. If no match is found, it falls back to semantic parsing to infer the intent and choose the reasoning modality.. Building on the task definitions established in existing graph understanding benchmarks Chen et al. (2024a); Yuan et al. (2025a); Cao et al. (2025), we propose a task classification framework that leverages the complementary strengths of different graph representation modalities. Specifically, graph tasks are categorized into two classes, each handled by a dedicated module:
Simple Property Task Examples: “What is the total number of nodes in this Graph?”; “Return the degree of node A.” Complex Local Task Examples: “Find the shortest path between nodes A and B.”; “Do nodes A and B share any third-order neighbors?”
Simple Property Tasks.
These tasks involve questioning either macroscopic graph statistics or microscopic, explicitly defined properties. Such information is efficiently and reliably obtained from textual graph descriptions, whereas visual representations often suffer from resolution limitations that obscure fine-grained details. Accordingly, the planning agent routes these tasks to the GraphRAG agent module, providing a high-level retrieval strategy for execution over $\mathcal{K}$.
Complex Local Tasks.
@@
These tasks require reasoning over fine-grained topological structures, such as finding the shortest path between nodes. For such tasks, a visual subgraph offers a concise and intuitive representation that is more effective for complex topological reasoning than verbose textual descriptions. Consequently, the planning agent delegates such tasks to the Visual Graph Thoughts agent, while decomposing the task into algorithm-guided substeps and constructing a high-level reasoning plan.
3.4 Visual Modality-based Graph Understanding
This module leverages visual-modality graph representations to handle complex local problems that require fine-grained topological understanding and multi-step logical inference.
To address the scalability issue of large graphs, we extract a task-relevant $k$-hop subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ with a maximum of $N_{max}$ nodes from $G$ in $\mathcal{K}$, centered on the key entities $\mathcal{E}$ (details are provided in Appendix B). The resulting subgraph is then transformed into a high-resolution visual representation $G_{\text{image}}=f_{\mathrm{viz}}(G^{\prime})$, which preserves the essential topological structure of $G^{\prime}$ and serves as a faithful basis for downstream reasoning.

### v2

# See or Say Graphs: Agent-Driven Scalable Graph Structure Understanding with Vision-Language Models

@@ Abstract
See or Say Graphs: Agent-Driven Scalable Graph Structure
Understanding with Vision-Language Models
@@ 1 Introduction
Recently, vision-language models (VLMs) have shown potential for general-purpose graph structure understanding Tang et al. (2024a); Kong et al. (2025); Ding et al. (2025b), offering new paradigms for solving real-world problems naturally represented as graphs. A central research focus of this emerging field is to enhance the foundational capability of VLMs to understand graph structures Ren et al. (2024), particularly for large-scale graphs, without relying on external tools such as code execution or software systems Ding et al. (2025a); Feng et al. (2025). To this end, existing studies have explored two primary graph input modalities for VLMs: textual graph descriptions [^1] Most early studies Cao et al. (2025); Guan et al. (2025); Peng et al. (2025) on text-modality-based graph understanding were developed on LLMs. As VLMs extend LLMs with visual capabilities, this paper presents a unified discussion from the perspective of graph structure understanding with VLMs., which convert structural information into natural language for indirect memory and reasoning over node properties and relations Jiang et al. (2023); Jin et al. (2024); Zhang et al. (2024a); and visual graph representations, which render nodes and edges as structured images to support direct structural perception Li et al. (2024c); Zhu et al. (2025b). More recent efforts further combine these modalities to improve structural understanding and complex reasoning on graphs Wei et al. (2024).
@@
Figure 1: Comparison of textual and visual modalities across graph structure understanding tasks
@@
However, despite these, current VLMs still lack a unified framework that systematically addresses the scalability challenges and the difficulties in coordinating textual and visual modalities inherent in graph structure understanding. Specifically:
@@
Scalability remains a fundamental limitation for both textual and visual graph inputs in VLMs. Given their limited input token capacities, textual graph descriptions are constrained by sequence length, making it difficult to encode the full structure of large graphs Chen et al. (2024a); Yuan et al. (2025a). Visual graph inputs, on the other hand, are restricted by image resolution, resulting in the loss of structural details as graph size increases Zhu et al. (2025b).
@@
In addition, VLMs face the lack of an effective mechanism to coordinate different input modalities for graph structure understanding, which prevents them from fully leveraging the distinct strengths and complementary advantages of textual and visual modalities. As illustrated in Figure 1, under the token length limits of VLMs, the text modality is more effective for tasks that require direct statistical access to explicit graph properties, such as counting nodes, measuring degrees, and assessing connectivity. In these cases, the information can be directly located and reasoned over with minimal effort. By contrast, the visual modality often performs worse because the limited resolution of global visual graphs blurs structural details, causing VLMs to misidentify relevant elements and leading to performance degradation.
@@
For tasks that require local graph structure reasoning [^2] Tasks based on global graph structure require full-graph access coupled with fine-grained structural reasoning, and can be viewed as a composition of global information retrieval and local structural-reasoning tasks., such as shortest-path computation or cycle detection, high-resolution local visual subgraphs offer a clear advantage for VLMs. They provide a “what-you-see-is-what-you-get” view of the graph topology and naturally support structural reasoning. In contrast, textual descriptions are often lengthy and lack structural clarity, leaving VLMs with only indirect language cues and thus little structural grounding for effective complex graph reasoning.
@@
Therefore, a natural conclusion is that fully leveraging VLMs for graph structure understanding requires addressing these two challenges. This entails, on the one hand, preserving as much task-relevant graph information as possible within limited input tokens, and on the other hand, assigning tasks according to modality strengths: letting the text modality efficiently address graph property tasks, while leveraging the visual modality to directly support local structure reasoning tasks.
@@
To this end, we propose GraphVista, an agent-driven unified framework designed to coordinate textual and visual modalities for robust graph understanding, ensuring the systematic robustness required to reliably tackle challenging graph tasks. For the scalability challenges, GraphVista is inspired by the retrieval-augmented generation (RAG) paradigm Lewis et al. (2020). It organizes graph information hierarchically according to structural importance and stores multi-granularity textual descriptions in a lightweight GraphRAG base. For a given graph structure understanding task, GraphVista retrieves only the relevant information before input, thereby compressing redundant context and preserving key reasoning elements. When necessary, it can also generate high-resolution visual subgraphs from the retrieved local information to support complex reasoning tasks.
@@
To enable effective coordination across modalities, GraphVista is built upon a planning agent that drives the overall workflow. It parses graph tasks, decomposes composite tasks if necessary, and routes them to the appropriate modality. The textual modality branch, enhanced with the RAG techniques, efficiently handles graph properties tasks by retrieving and reasoning over concise textual descriptions. In contrast, the visual modality branch employs the Visual Graph Thoughts Agent, which constructs a multimodal chain-of-thought called Visual Graph Thoughts over high-resolution visual subgraphs, enabling VLMs to perform stepwise visual reasoning grounded in explicit structural evidence for visual tasks.
The main contributions are as follows:
@@
    We present the first systematic analysis of the scalability challenges and coordination gaps that hinder VLMs in graph structure understanding, providing both conceptual insights and empirical evidence.
@@
    We propose GraphVista, the first unified framework for graph structure understanding with VLMs, designed around two objectives: improving scalability and enabling coordination across modalities. To this end, GraphVista integrates a set of innovative techniques, such as a hierarchical GraphRAG base, a planning agent for task routing, and a multimodal graph thoughts agent for complex reasoning.
@@
    We present Grena, the first large-scale graph benchmark designed to support step-level evaluation of multimodal graph understanding.
@@
    Extensive experiments across diverse graph tasks demonstrate the effectiveness of GraphVista, highlighting its ability to scale to large graphs and to exploit the complementary strengths of textual and visual modalities.
@@ 2 Related Work
2.1 Text Modality-based Graph Structure Understanding
Most existing studies convert graphs into textual descriptions to enable VLM-based graph structure understanding, primarily focusing on benchmark construction and capability analysis. Recent benchmarks evaluate VLMs on basic graph element comprehension Wang et al. (2023); Guo et al. (2023); Wu et al. (2024) and graph-theoretic reasoning tasks Chen et al. (2024a); Li et al. (2024b); Tang et al. (2024b); Luo et al. (2024); Yuan et al. (2025a), indicating that the research in this direction remains in an early exploratory phase. Methodologically, Ge et al. (2025); Wang et al. (2025) studied the influence of graph description order, while Cao et al. (2025) identified the “Lost in the Middle” issue in graph sequence. Building upon these findings, Guan et al. (2025) analyzed VLM attention patterns over graph data, and Peng et al. (2025); Zhang et al. (2025) explored task transferability to broader reasoning domains. However, these strategies still face scalability bottlenecks and struggle with complex reasoning tasks.
@@ 2.1 Text Modality-based Graph Structure Understanding
2.2 Visual Modality-based Graph Structure Understanding
Visual modality-based graph structure understanding is still in its infancy. Li et al. (2024c), Zhu et al. (2025b), and Zhao et al. (2025) benchmark VLMs on graph perception and reasoning, revealing difficulties in capturing structural information. Wei et al. (2024) further introduces a hybrid multimodal representation directly compatible with VLMs. However, these studies remain empirical and fail to address the scalability, structural understanding, and reasoning limitations inherent to vision-only representations.
@@ 3.1 Overview
We introduce GraphVista, an agent-driven framework designed to address the challenges of scalability and modality coordination in graph structure understanding [^3] Aligned with prior studies Wang et al. (2023); Chen et al. (2024a); Yuan et al. (2025a), our work focuses on pure graph structure understanding. The processing of semantic information is orthogonal to our method and can be seamlessly incorporated into GraphVista due to its modular design. , as shown in Figure 2.
GraphVista is coordinated by a planning agent that directs the entire workflow. Initially, the planning agent constructs a lightweight Hierarchical GraphRAG Base $\mathcal{K}$ by partitioning the graph based on topological centrality. Upon receiving a question, the agent parses the request to identify the task type and key entities, decomposes composite tasks if necessary, and routes them to the appropriate reasoning modality. For text-modality tasks, the text modality-based branch is activated, employing GraphRAG-based retrieval to efficiently extract relevant context from $\mathcal{K}$. For visual-modality tasks, the task is delegated to the visual modality-based branch. In this branch, the Visual Graph Thoughts Agent extracts task-relevant visual subgraphs from $\mathcal{K}$ and performs iterative multimodal reasoning grounded in explicit visual evidence.
The remainder of this section outlines our proposed framework. We begin with the construction of the hierarchical base $\mathcal{K}$ in §3.2, followed by the planning agent-driven task routing mechanism in §3.3. Finally, we elaborate on the graph understanding modules specific to the visual and textual modalities in §3.4 and §3.5, respectively.
@@ 3.2 Hierarchical GraphRAG Base
Limited context windows often prevent VLMs from processing complete graphs, leading to the loss of key structural details. To address this, we propose constructing a local knowledge base $\mathcal{K}$. Diverging from conventional RAG pipelines Guo et al. (2025); Jin et al. (2025) that rely on external knowledge sources, our approach constructs a compact multi-level structural representation directly from the graph and stores it in a lightweight GraphRAG Base, enabling scalable reasoning without external dependencies [^4] For small graphs ($|V|\leq 15$), we skip the storage phase and feed the entire graph directly into the VLMs..
Specifically, we partition the graph $G=(V,E)$ with $|V|$ nodes into three tiers based on the topological centrality of its nodes. Each tier uses a different storage granularity, significantly reducing storage needs while preserving the structural details essential for downstream tasks.
@@
This tier comprises nodes with the strongest structural influence, acting as hubs that govern information flow and global graph properties. We employ PageRank, which captures both connectivity and reachability, to identify the top $K_{1}\%$ of nodes as core nodes. For each, we preserve the 2-hop neighborhood [^5] We limit the storage scope to a 2-hop radius to optimize the trade-off between structural coverage and computational cost. Extending to 3-hops (or beyond) leads to “neighbor explosion” Barabási and Albert (1999); Grabow et al. (2015), yet remote nodes contribute little to local structural reasoning Li et al. (2018); Liben-Nowell and Kleinberg (2003). We adopt this generalized strategy to ensure broad applicability; advanced GraphRAG storage optimizations are orthogonal to our contribution and can be seamlessly integrated. $\mathcal{N}_{2}(v)$ to explicitly retain fine-grained local topology.
@@
Situated along inter-community paths, these nodes are critical for maintaining global connectivity. We select them using Betweenness Centrality, prioritizing nodes that frequently bridge shortest paths between pairs. The subsequent $K_{2}\%$ of nodes are retained; however, to balance representational fidelity with storage efficiency, we store only the 1-hop neighborhood $\mathcal{N}_{1}(v)$ for each backbone node.
@@ 3.3 Planning Agent-based Task Routing
This module employs a planning agent to allocate each task to the optimal reasoning modality, either textual or visual.
Task Parsing.
Given a natural language question $Q$ over $G$, the planning agent employs semantic parsing to determine the task type $T$ and extract relevant entities $\mathcal{E}$:
@@
, where $f_{\text{parse}}$ utilizes few-shot prompting [^7] The planning agent first matches the question against predefined task templates; if unmatched, it falls back to semantic parsing to infer the intent and choose the reasoning modality. Implementation details and a quantitative analysis of task classification error rates are provided in Appendix H. to infer task semantics, producing a task category $T$ (e.g., neighbor identification) and a set of entity pairs $\mathcal{E}$. To ensure consistent downstream processing, queries involving a single entity (e.g., node degree) are normalized to $(v_{i},\emptyset)$ using a null placeholder.
Task Routing and Execution Planning.
After parsing the task, the planning agent routes the problem to the appropriate processing module based on the task type $T$ and formulates a detailed execution plan. Integrating task designs from existing graph understanding benchmarks, the planning agent routes tasks based on their dependence on structural reasoning, thereby aligning each task with the respective strengths of text and vision modalities [^8] Composite tasks that can be decomposed into single-modality operations (e.g., determining which of two nodes shares more common neighbors with a given node) can be handled by the planning agent through sequential execution of the corresponding subtasks.. Specifically, tasks that do not require structural reasoning are handled by the text modality, tasks requiring local structural reasoning are assigned to the visual modality, and tasks involving global graph structure are decomposed into stepwise text- and vision-based subtasks.
Text-Modality Task Examples: “What is the total number of nodes in this graph?”; “What is the degree of node 15?” Visual-Modality Task Examples: “Find the shortest path between nodes A and B.”; “Whether node A is part of any triangle.” Modality-Collaborative Task Examples: “Calculate the diameter of the graph.” Decomposition: (i) Identify peripheral nodes using text; (ii) Estimate pairwise distances through visual subgraph reasoning.”
Text-Modality Tasks.
These tasks do not require structural reasoning over the graph. Instead, they rely on direct access to individual nodes or edges for basic statistical retrieval, such as counting the total number of nodes or querying the degree of a given node. Accordingly, the planning agent directs these tasks to the GraphRAG agent module, providing a retrieval strategy for execution over $\mathcal{K}$.
Visual-Modality Tasks.
@@
These tasks involve reasoning over local graph structures rather than simple retrieval. They focus on topological relationships within a limited subgraph (e.g., shortest-path tasks). In such cases, visual subgraph representations provide a clearer and more efficient form for structural reasoning than textual descriptions. Consequently, the planning agent delegates such tasks to the visual graph thoughts agent, while decomposing the task into algorithm-guided substeps and constructing a high-level reasoning plan.
Modality-Collaborative Tasks.
These tasks rely on global graph structure understanding that combines full-graph access with fine-grained structural analysis, which is often beyond a single modality. They are therefore formulated as a sequence of steps that alternate between global information access and local structural reasoning. The planning agent decomposes the task into sequential text- and visual-modality sub-tasks.
3.4 Visual Modality-based Graph Structure Understanding
This module leverages visual-modality graph representations to handle visual-modality tasks that require fine-grained topological understanding and multi-step logical inference. To address scalability, we extract a task-relevant $k$-hop subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ centered on key entities $\mathcal{E}$, with at most $N_{\max}$ nodes (as detailed in Appendices B and C). $G^{\prime}$ is visualized as $G_{\text{image}}=f_{\mathrm{viz}}(G^{\prime})$ via a task-driven strategy that adapts layouts to the task and simplifies dense subgraphs for visual clarity, ensuring that essential structural information remains perceivable to the VLM.